Please help with all three parts
Three point charges are at z 1.2m F Value
Answer #1

q1 = 18 μC, x1 = 0, q2 = 32 μC, x2 = 0.60 m, q3 = -10 μC, x3 = 1.2 m

A) force on q1 = -kq1q2/(x2 - x1)^2 + kq1|q3|/(x3 - x1)^2 = -13.1 N
q1 = (-9*10^9*18*10^-6*32*10^-6)/(0.60-0)^2 + (9*10^9*18*10^-6*10*10^-6)/(1.2-0)^2
q1 = 13.27 N

B) force on q2 = kq1q2/(x2 - x1)^2 + kq2|q3|/(x3 - x2)^2 = 22.4 N

C) force on q3 = -kq1|q3|/(x3 - x1)^2 - kq2|q3|/(x3 - x2)^2 = -9.12 N